3. Check whether 7 + 3x U1 a factor of + 71

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
If 7 + 3% is a fact.or of 3x3 + 7x, then p(-i) = 0 ( 7) (-7)3 .. p -3 = 3 x 3 + 7 x (-37) -343 49 -343 49 = 3 x 27 - 3 = 9- - 3 -343-147 -490 = 9 = 9- • 0• As remainder is not zero, hence (7 + 3%) is not a factor of 3.%3 + 7x.